OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)).

OpenStudy (shamil98):

f(g(x)) plug in the equation of g(x) into f(x).

OpenStudy (shamil98):

f(x) = 9x - 2 g(x) = -x+ 3 9(-x+3) - 2 = f(g(x)) simplify it.